Find the general integral of the linear p.d.e



1) Px(x+y) =qy(x+y) - (x-y) (2x+2y+z)


2) x(x^2+3y^2) p - y(3x^2+y^2) q=2z(y^2-x^2)


3) (y+zx) p - (x+yz) q=x^2-y^2

1
Expert's answer
2021-12-07T13:09:22-0500

1)

"\frac{dx}{x(x+y)}=-\frac{dy}{y(x+y)}=-\frac{dz}{(x-y) (2x+2y+z)}"


"dx\/x+dy\/y=0"

"lnx+lny=lnc_1"

"xy=c_1"


"\frac{dx+dy}{x+y}=\frac{-dz}{2(x+y)+z}"


let "u=x+y,z=tu" , then:

"z'=t'u+t=-\frac{2u+tu}{u}=-2-t"


"-\frac{dt}{2(t+1)}=du/u"


"-ln(t+1)\/2=lnu+lnc"

"\frac{1}{\sqrt{t+1}}=cu"


"u\sqrt{t+1}=(x+y)\sqrt{\frac{z}{x+y}+1}=c_2"


"F(c_1,c_2)=F(xy,(x+y)\sqrt{\frac{z}{x+y}+1})=0"


2)

"\frac{dx}{ x(x^2+3y^2)}=-\frac{dy}{y(3x^2+y^2)}=\frac{dz}{2z(y^2-x^2)}"


"dx\/x+dy\/y=dz\/z"

"lnx+lny=lnz+lnc_1"

"xy\/z=c_1"


"-\frac{y(3x^2+y^2)}{ x(x^2+3y^2)}=\frac{dy}{dx}"


"y=tx,y'=t'x+t"


"t'x+t=-\frac{t(3+t^2)}{1+3t^2}"


"t'x=-\frac{3t+t^3+t+3t^3}{1+3t^2}"


"\frac{1+3t^2}{t^3+t}dt=-\frac{4dx}{x}"


"ln(t^3+t)=-4lnx+lnc_2"

"x(t^3+t)=x((y\/x)^3+y\/x)=c_2"

"y^3\/x^2+y=c_2"


"F(c_1,c_2)=F(xy\/z,y^3\/x^2+y )=0"


3)

"\frac{dx}{y+zx}=-\frac{dy}{x+yz}=\frac{dz}{x^2-y^2}"


"ydx+xdy+dz=0"

"xy+z^2\/2=c_1"


"\frac{xdx+ydy}{z(x^2-y^2)}=\frac{dz}{x^2-y^2}"


"x^2+y^2-z^2=c_2"


"F(c_1,c_2)=F(xy+z^2\/2, x^2+y^2-z^2)=0"